### v1

@@ 4 Conclusion
This paper has developed a strategic framework for smart grids with many agents using mean field game methodology. By analyzing both cooperative and non-cooperative scenarios under sticky prices and finite time horizons, we provide approaches to minimize expected losses through social cost reduction and approximate Nash equilibrium, respectively. Numerical experiments demonstrate the effectiveness of these approximations, offering insights for managing large-scale smart grid systems.
@@ Appendix A Proofs of Auxiliary Lemmas
For any $i$, we can consider all agents except for agent $i$ as a large population which consists of N1N-1 agents. Thus, by applying our previous calculations, we obtain
@@
To proceed, we require an exact expression of $\hat{P}$. We directly solve from (2.2)
@@
$\leq O(1)+(\alpha T)^{2}\mathbb{E}\left[\sup_{0\leq t\leq T}|{\hat{Q}_{t}}|^{2}\right]$
@@
where we have applied (A). This shows that
$\mathbb{E}\left[\int_{0}^{T}\hat{P}_{t}^{2}dt\right]\leq T\mathbb{E}\left[\sup_{0\leq t\leq T}|{\hat{P}_{t}}|^{2}\right]\leq O(1).$
Moreover, (A) gives
@@
This proves that
@@
Let $K^{\prime}$ be an $O(1)$ number that exceeds $K$ and all the $O(1)$ numbers above, and we complete the proof of Lemma 2.1. ∎
Here are the proofs of Lemma 2.2 and Lemma 2.3.
@@
Note that $v\in\mathcal{A}$ implies
@@
Consequently, we adopt $\int_{0}^{T}v_{t}^{2}dt<\infty$ in what follows, taking into account that events within a zero-probability set do not affect the expectation. In this setting, the integral as shown in Lemma 2.2 exists and is finite. To present the proof clearly, we start by introducing some notation. We set
@@
Our objective is to apply Positive Real Lemma (Boyd et al. (1994)) to the system. To proceed, we denote the LMI for the system by
@@
We explore the conditions under which the LMI is semi-negative definite. By plugging in the exact value of $a^{*}$ and $b^{*}$, we discover that
@@
It is assumed that $N\geq\frac{\sqrt{2}+1}{2{\epsilon_{2}}^{*}}$, so the LMI is semi-negative definite. Positive Real Lemma implies
@@
By plugging in the exact form of $C_{1}$ and $D_{1}$ and taking the expectation, we immediately obtain that Lemma 2.2 holds true.
@@
Fix positive numbers $\epsilon_{1},\epsilon_{2},\epsilon_{3},\epsilon_{4}$ and $\epsilon_{5}$, all independent of $N$, so that
@@
Our first step relies on (2.2), which gives
@@
To evaluate how $P$ changes in response to $v$, we express $P$ as the sum of two components, one dependent on $v$ and the other independent of $v$. As shown below, we set
@@
We naturally have $P_{t}=P_{t}^{\prime}+P_{t}^{\prime\prime}$. The above expression for $P^{\prime\prime}_{t}$ is equivalently characterized by the dynamics $dP_{t}^{\prime\prime}=-\alpha P_{t}^{\prime\prime}-\frac{\alpha}{N}v_{t}^{i}$, $P^{\prime\prime}_{0}=0$. $X_{t}^{i}$ and $P_{t}^{\prime\prime}$ are both processes induced by $v^{i}$, and the coefficients are similar to those in Lemma 2.2. However, to apply this lemma requires that the initial values for $X_{t}^{i}$ and $P_{t}^{\prime\prime}$ equal zero, which is not possible to be fulfilled in the general case. Therefore, we need to take a detour. We construct the process below,
@@
Note that the dynamics and initial value of $(\dot{X}_{t}^{i},\ P^{\prime\prime}_{t})^{\top}$ and that in Lemma 2.2 are identical. We then substitute $({X}_{t}^{*},\ P_{t}^{*})^{\top}$ with $(\dot{X}_{t}^{i},\ P^{\prime\prime}_{t})^{\top}$ in Lemma 2.2. It follows that, if $N>\frac{\sqrt{2}+1}{2{\epsilon_{2}}}$, we have
@@
Combining $P_{t}=P_{t}^{\prime}+P_{t}^{\prime\prime}$ and (A.2), we obtain an inequality for $J_{i}$ satisfied by $\forall N>\frac{\sqrt{2}+1}{2{\epsilon_{2}}}$,
@@
From Lemma 2.1, direct calculation and the Cauchy-Schwartz inequality yield that
@@
$\quad+2\mathbb{E}\left[\sup_{t\in[0,T]}\left\{\left(\int_{0}^{t}\frac{\alpha}{N}e^{-\alpha(t-s)}\bigg{(}\sum_{k\neq i}\hat{v}_{s}^{k}\bigg{)}ds\right)^{2}\right\}\right]$
$\leq O(1)+(\alpha T)^{2}O(1)$
$\leq O(1).$
This implies that
@@
The fundamental inequality gives $|P^{\prime}_{t}v_{t}^{i}|\leq\frac{\epsilon_{5}}{2}(v_{t}^{i})^{2}+\frac{1}{2\epsilon_{5}}(P^{\prime}_{t})^{2}$, which leads to
@@
This simplifies (A) into the following inequality, which holds true once $N$ is larger than $\frac{\sqrt{2}+1}{2{\epsilon_{2}}}$,
@@
We want to cover the possible negative term $\epsilon_{1}\dot{X}_{t}^{i}v_{t}^{i}$ with a sum of several multiples of $(\dot{X}_{t}^{i})^{2}$ and $(v_{t}^{i})^{2}$. However, we only have $X_{t}^{i}$ in this inequality. Thus, we reduce $\mathbb{E}[(X_{t}^{i}-\kappa)^{2}]$ into an expression of $\dot{X}_{t}^{i}$. We obtain from $(X_{t}^{i}-\kappa)-(x_{0}^{i}+\sigma_{i}W_{t}^{i}-\kappa)=\dot{X}_{t}^{i}$,
@@
$\leq 2\mathbb{E}[(X_{t}^{i}-\kappa)]^{2}+4(x_{0}^{i}-\kappa)^{2}+4\sigma_{i}^{2}T.$
This gives
$\mathbb{E}\left[(X_{t}^{i}-\kappa)^{2}\right]\geq$ $\frac{1}{2}\mathbb{E}\left[(\dot{X}_{t}^{i})^{2}\right]+O(1).$
Then,
@@
$\quad+\mathbb{E}\left[\int_{0}^{T}\left(-\epsilon_{1}\dot{X}_{t}^{i}v_{t}^{i}+\epsilon_{4}(v_{t}^{i})^{2}+\frac{\epsilon_{3}}{2}(\dot{X}_{t}^{i})^{2}\right)dt\right]+O(1)\$
is satisfied for $\forall N>\frac{\sqrt{2}+1}{2{\epsilon_{2}}}=\frac{4(\sqrt{2}+1)}{c}$. From the restrictions on $\epsilon_{1},\epsilon_{2},\epsilon_{3},\epsilon_{4}$ and $\epsilon_{5}$, the contents of the first two parentheses provide the desired leading term. Observe that, by the restriction $\epsilon_{1}^{2}=\epsilon_{3}\epsilon_{4}$, fundamental inequality implies $\epsilon_{1}\dot{X}_{t}^{i}v_{t}^{i}+\epsilon_{4}(v_{t}^{i})^{2}+\frac{\epsilon_{3}}{2}(\dot{X}_{t}^{i})^{2}\geq 0$. Therefore, the proof is complete.
The proofs of Lemma 2.4 and Lemma 2.5 follow.
@@
For $v\in\mathcal{A}$, we adopt the condition $\int_{0}^{T}v_{t}^{2}dt<\infty$ in the analysis below, as in the proof of Lemma 2.2. We introduce the notion
$A_{2}=\begin{pmatrix}0&0\\ 0&-\alpha\end{pmatrix}\preceq O,\quad B_{2}=\begin{pmatrix}1\\ -\alpha\end{pmatrix},\quad C_{2}=\begin{pmatrix}\alpha&1\end{pmatrix},\quad D_{2}=1.$
Recalling Lemma 2.4’s statement, we introduce the output function
$\tilde{y}_{t}:=v_{t}+\alpha\tilde{X}_{t}+\tilde{P}_{t}=C_{2}\tilde{x}_{t}+D_{2}.$
Then, the process simplifies to
@@
We state that $L_{2}\preceq 0$, for it is equivalent to
@@
This is obvious. From the Positive Real Lemma, we conclude that
$\int_{0}^{T}v_{t}\tilde{y}_{t}dt=\int_{0}^{T}v_{t}(v_{t}+\alpha\tilde{X}_{t}+\tilde{P}_{t})dt\geq 0.$
We complete the proof by taking the expectation. ∎
@@
Making use of Assumption $(\boldsymbol{A_{4}})$, we fix $\epsilon,\epsilon^{\prime}>0$ small enough, $A\in(0,\frac{\eta}{2})$, and $h\in(0,1)$ so that
@@
In order to apply Lemma 2.4, which requires the starting value to be 0, we introduce three processes $P_{t}^{\dagger}$, $\overline{X}_{t}$ and $\overline{X}_{t}^{\prime}$, whose dynamics satisfy
@@
Note that $\overline{X}_{t}$ is the average of $X_{t}^{i}$, so it contains a Brownian motion term and does not start at zero. The other two processes start at zero, and do not contain a Brownian motion term in their dynamics. From their definition, we get
$\overline{X}_{t}=\bar{x}_{0}^{N}+\overline{X}_{t}^{\prime}+\sigma B_{t}.$
A simple calculation shows
@@
The dynamics and initial value of $(\overline{X}_{t}^{\prime},\ P_{t}^{\dagger})^{\top}$ are the same as that of the process displayed in Lemma 2.4. Therefore, Lemma 2.4 implies that
@@
(A.6) involves $P_{t}^{\dagger}$ and $\overline{X}_{t}^{\prime}$, which do not appear in $J_{soc}$. Therefore, we aim to reduce $J_{soc}$ into an expression where $\frac{\eta}{2}(X_{t}^{i}-\kappa)^{2}$ and part of $\frac{c}{2}v_{t}^{i}$ are substituted by expressions of $P_{t}^{\dagger}$, $Q_{t}$ and $\overline{X}_{t}^{\prime}$. Our first step relies on the convexity of $L$, which indicates that $\sum_{i=1}^{N}\frac{\eta}{2}(X_{t}^{i}-\kappa)^{2}$ is not smaller than $\frac{\eta N}{2}(\overline{X}_{t}-\kappa)^{2}$, and $\sum_{i=1}^{N}(\frac{c}{2}-\epsilon^{\prime})(v_{t}^{i})^{2}$ is not smaller than $\sum_{i=1}^{N}(\frac{c}{2}-\epsilon^{\prime})Q_{t}^{2}$. In light of this, we write
@@
Our second step relies on the exact expression of $P$. To separate the control-independent components, we introduce the following definition
$C(t)=e^{-\alpha t}p_{0}+\int_{0}^{t}e^{-\alpha(t-s)}\alpha\beta ds,\quad t\in[0,T].$
This function is obviously equipped with a $\|\cdot\|_{\infty}$ bound over $[0,T]$. The exact form of $P_{t}$, along with (A.5), shows that
@@
Our third step relies on the relationship between $\overline{X}_{t}$ and $\overline{X}_{t}^{\prime}$. We have
@@
Plug (A), (A.9) into (A.7), and rearrange the terms to get another inequality for $J_{soc}$
@@
$\quad+\frac{1}{N}\mathbb{E}\left[\int_{0}^{T}\left(C(t)Q_{t}+\epsilon Q_{t}^{2}+\epsilon^{\prime}{(v_{t}^{i})}^{2}\right)dt\right]+O(1).$
We have derived from Lemma 2.4 that $\mathbb{E}\left[\int_{0}^{T}(P_{t}^{\dagger}Q_{t}+Q_{t}^{2}+\alpha Q_{t}\overline{X}_{t}^{\prime})dt\right]\geq 0$, and (A.4), together with the fundamental inequality, shows that $(Ah\overline{X}_{t}^{\prime 2}+(\frac{c}{2}-\epsilon^{\prime}-1-\epsilon)Q_{t}^{2}-\alpha Q_{t}\overline{X}_{t}^{\prime})$ stays non-negative. These observations lead us to further reduce the inequality for $J_{soc}$ into
@@
Note that $(\frac{\eta}{2}h-Ah)\overline{X}_{t}^{\prime 2}+\frac{\eta h}{2}\big{(}({\bar{x}_{0}^{N}})^{2}+2\overline{X}_{t}^{\prime}\bar{x}_{0}^{N}\big{)}-\eta\kappa\overline{X}_{t}^{\prime}$ and $\frac{\epsilon}{2}Q_{t}^{2}-\|C(t)\|_{\infty}|Q_{t}|$ have an $O(1)$ lower bound, so
$J_{soc}(\boldsymbol{v})\geq\frac{1}{N}\mathbb{E}\left[\sum_{i=1}^{N}\left(\int_{0}^{T}\left(\frac{\epsilon}{2}Q_{t}^{2}+\epsilon^{\prime}(v_{t}^{i})^{2}\right)dt\right)\right]+O(1).$
The proof of Lemma 2.5 is complete.

### v2

@@ 6 Conclusion
This paper has developed a strategic framework for smart grids with many agents using mean field game methodology. Under sticky prices and finite time horizons, we analyze both the cooperative and non-cooperative settings. In the cooperative case, we derive strategies that minimize the expected social cost, while in the non-cooperative case, we construct approximate Nash equilibrium. Numerical experiments demonstrate the effectiveness of these approximations, offering insights for managing large-scale smart grid systems.
@@ Appendix A Proofs of Auxiliary Lemmas
For any $i$, we may treat all other agents as a large population of which consists of N1N-1 agents. Applying the calculations from the previous calculations, we obtain
@@
To proceed, we require an exact expression of $\hat{P}$. Solving directly from (2.2) yields
@@
$\leq O(1)+2(\alpha T)^{2}\mathbb{E}\left[\sup_{0\leq t\leq T}|{\hat{Q}_{t}}|^{2}\right]$
@@
where we have used the bounds in (A). This shows that
$\mathbb{E}\left[\int_{0}^{T}(\hat{P}_{t})^{2}dt\right]\leq T\mathbb{E}\left[\sup_{0\leq t\leq T}|{\hat{P}_{t}}|^{2}\right]\leq O(1).$
Moreover, (A) implies that
@@
This establishes that
@@
Let $K^{\prime}$ be an $O(1)$ number that exceeds $K$ and all the $O(1)$ numbers above. With this, the proof of Lemma 3.1 is complete. ∎
We now turn to the proofs of Lemma 3.2 and Lemma 3.3.
@@
Note that $v\in\mathcal{A}_{i}$ implies
@@
Consequently, we restrict attention to events for which the above integral is finite. Under such events, the integral in Lemma 3.2 exists and is finite. To present the proof clearly, we introduce the following notation:
@@
Our objective is to apply Positive Real Lemma (Section 2.7.2, Boyd et al. (1994)) to the system. To proceed, we denote the LMI for the system by
@@
We investigate the conditions under which the LMI is semi-negative definite. By plugging in the exact value of $a^{*}$ and $b^{*}$, we discover that
@@
Since we assume $N\geq\frac{\sqrt{2}+1}{2{\epsilon_{2}}^{*}}$, the LMI is semi-negative definite. Positive Real Lemma implies
@@
Substituting the explicit forms of $C_{1}$ and $D_{1}$ and taking the expectation, we immediately obtain that Lemma 3.2 holds.
@@
Fix positive numbers $\epsilon_{1},\epsilon_{2},\epsilon_{3},\epsilon_{4}$ and $\epsilon_{5}$ such that
@@
Our first step uses (2.2), which gives
@@
To evaluate how $P$ changes in response to $v$, we decompose it into two components, one dependent on $v$ and the other independent of $v$. As shown below, we set
@@
We naturally have $P_{t}=P_{t}^{\prime}+P_{t}^{\prime\prime}$. The above expression for $P^{\prime\prime}_{t}$ is equivalently characterized by the dynamics $dP_{t}^{\prime\prime}=-\alpha P_{t}^{\prime\prime}-\frac{\alpha}{N}v_{t}^{i}$, $P^{\prime\prime}_{0}=0$. Both processes $X_{t}^{i}$ and $P_{t}^{\prime\prime}$ are driven by $v^{i}$, and their coefficients match those in Lemma 3.2. However, applying the lemma directly requires that both initial conditions be zero, which does not hold in general. Thus, we take a detour and construct the process
@@
Note that the dynamics and initial value of $(\dot{X}_{t}^{i},\ P^{\prime\prime}_{t})^{\top}$ coincide with those in Lemma 3.2. Therefore, we substitute $({X}_{t}^{*},\ P_{t}^{*})^{\top}$ with $(\dot{X}_{t}^{i},\ P^{\prime\prime}_{t})^{\top}$ in Lemma 3.2, and obtain that if $N>\frac{\sqrt{2}+1}{2{\epsilon_{2}}}$, then
@@
Combining $P_{t}=P_{t}^{\prime}+P_{t}^{\prime\prime}$ and (A.2), we have, for all $N>\frac{\sqrt{2}+1}{2{\epsilon_{2}}}$,
@@
From Lemma 3.1, direct calculations and the Cauchy–Schwarz inequality yield
@@
$\quad+2\mathbb{E}\left[\sup_{t\in[0,T]}\left\{\left(\frac{\alpha}{N}\int_{0}^{t}e^{-\alpha(t-s)}\bigg(\sum_{k\neq i}\hat{v}_{s}^{k}\bigg)ds\right)^{2}\right\}\right]$
$\leq O(1),$
and hence
@@
The fundamental inequality gives $|P^{\prime}_{t}v_{t}^{i}|\leq\frac{\epsilon_{5}}{2}(v_{t}^{i})^{2}+\frac{1}{2\epsilon_{5}}(P^{\prime}_{t})^{2}$, leading to
@@
This simplifies (A) into the following inequality, which holds once $N$ is larger than $\frac{\sqrt{2}+1}{2{\epsilon_{2}}}$,
@@
We now control the possible negative cross-term $\epsilon_{1}\dot{X}_{t}^{i}v_{t}^{i}$ using suitable multiples of $(\dot{X}_{t}^{i})^{2}$ and $(v_{t}^{i})^{2}$. Since the inequality contains $X_{t}^{i}$, we rewrite $\mathbb{E}[(X_{t}^{i}-\kappa)^{2}]$ in terms of $\dot{X}_{t}^{i}$. From $(X_{t}^{i}-\kappa)-(x_{0}^{i}+\sigma_{i}W_{t}^{i}-\kappa)=\dot{X}_{t}^{i}$, we get
@@
$\leq 2\mathbb{E}[(X_{t}^{i}-\kappa)]^{2}+4(x_{0}^{i}-\kappa)^{2}+4\sigma_{i}^{2}T,$
and hence
$\mathbb{E}\left[(X_{t}^{i}-\kappa)^{2}\right]\geq$ $\frac{1}{2}\mathbb{E}\left[(\dot{X}_{t}^{i})^{2}\right]+O(1),$
and, for $\forall N>\frac{\sqrt{2}+1}{2{\epsilon_{2}}}=\frac{4(\sqrt{2}+1)}{c}$
@@
$\quad+\mathbb{E}\left[\int_{0}^{T}\left(-\epsilon_{1}\dot{X}_{t}^{i}v_{t}^{i}+\epsilon_{4}(v_{t}^{i})^{2}+\frac{\epsilon_{3}}{2}(\dot{X}_{t}^{i})^{2}\right)dt\right]+O(1),$
$\geq\mathbb{E}\Bigg[\int_{0}^{T}\bigg(\left(\frac{\eta}{2}-\epsilon_{3}\right)(X_{t}^{i}-\kappa)^{2}+\left(\frac{c}{2}-\epsilon_{2}-\epsilon_{4}-\epsilon_{5}\right)(v_{t}^{i})^{2}\bigg)dt+O(1),\$
where the last step relies on the restriction $\epsilon_{1}^{2}=\epsilon_{3}\epsilon_{4}$. Furthermore, applying the restrictions on $\epsilon_{1},\epsilon_{2},\epsilon_{3},\epsilon_{4}$ and $\epsilon_{5}$ yield the desired result. Therefore, the proof is complete.
The proofs of Lemma 4.1 and Lemma 4.2 are presented below.
@@
For $v\in\mathcal{A}_{i}$, we impose the condition $\int_{0}^{T}v_{t}^{2}dt<\infty$, as in the proof of Lemma 3.2. We introduce the notion
$A_{2}=\begin{pmatrix}0&0\\ 0&-\alpha\end{pmatrix}\preceq O,\quad B_{2}=\begin{pmatrix}1\\ -\alpha\end{pmatrix},\quad C_{2}=\begin{pmatrix}\alpha&1\end{pmatrix},\quad D_{2}=1,$
and the output function
$\tilde{y}_{t}:=v_{t}+\alpha\tilde{X}_{t}+\tilde{P}_{t}=C_{2}\tilde{x}_{t}+D_{2},$
which allows us to rewrite the system as
@@
We state that $L_{2}\preceq 0$, which is equivalent to
@@
This is immediate. By the Positive Real Lemma, we conclude that
$\int_{0}^{T}v_{t}\tilde{y}_{t}dt=\int_{0}^{T}v_{t}(v_{t}+\alpha\tilde{X}_{t}+\tilde{P}_{t})dt\geq 0,$
and the proof follows upon taking expectations. ∎
@@
To proceed with the proof of Lemma 4.2, we make use of Assumption $(\boldsymbol{A_{4}})$. We fix $\epsilon,\epsilon^{\prime}>0$ small enough, along with $A\in(0,\frac{\eta}{2})$ and $h\in(0,1)$, such that
@@
To apply Lemma 4.1, which requires zero initial conditions, we introduce three processes $P_{t}^{\dagger}$, $\overline{X}_{t}$ and $\overline{X}_{t}^{\prime}$, whose dynamics satisfy
@@
$\overline{X}_{t}$, the average of $X_{t}^{i}$, contains a Brownian motion term and does not start from zero. The other two processes start at zero and have no Brownian components. From their definitions, we obtain
$\overline{X}_{t}=\bar{x}_{0}^{N}+\overline{X}_{t}^{\prime}+\sigma B_{t},$
and a direct computation shows that
@@
The dynamics and initial value of $(\overline{X}_{t}^{\prime},\ P_{t}^{\dagger})^{\top}$ coincide with those of the process in Lemma 4.1. Therefore, Lemma 4.1 implies that
@@
Since (A.6) involves $P_{t}^{\dagger}$ and $\overline{X}_{t}^{\prime}$, which do not appear in $J_{soc}$, we aim to express $J_{soc}$ in a form where $\frac{\eta}{2}(X_{t}^{i}-\kappa)^{2}$ and part of $\frac{c}{2}v_{t}^{i}$ are replaced by terms of $P_{t}^{\dagger}$, $Q_{t}$ and $\overline{X}_{t}^{\prime}$. Our first step uses the convexity of $L$, which indicates that
@@
Our second step is based on separating the control-independent part of the price. We introduce the function
$C(t)=e^{-\alpha t}p_{0}+\int_{0}^{t}e^{-\alpha(t-s)}\alpha\beta ds,\quad\forall t\in[0,T],$
which is uniformly bounded on $[0,T]$. Using the exact form of $P_{t}$, we have
@@
Our third step concerns the relationship between $\overline{X}_{t}$ and $\overline{X}_{t}^{\prime}$.
@@
Substituting (A) and (A.9) into (A.7) and rearranging terms yield
@@
$\quad+\frac{1}{N}\mathbb{E}\left[\int_{0}^{T}\left(C(t)Q_{t}+\epsilon Q_{t}^{2}+\epsilon^{\prime}{(v_{t}^{i})}^{2}\right)dt\right]+O(1).$ (A.10)
We have shown via Lemma 4.1 that $\mathbb{E}\left[\int_{0}^{T}(P_{t}^{\dagger}Q_{t}+Q_{t}^{2}+\alpha Q_{t}\overline{X}_{t}^{\prime})dt\right]\geq 0$, and by (A.4) and the fundamental inequality, $(Ah\overline{X}_{t}^{\prime 2}+(\frac{c}{2}-\epsilon^{\prime}-1-\epsilon)Q_{t}^{2}-\alpha Q_{t}\overline{X}_{t}^{\prime})$ stays non-negative. Hence we reduce (A) into
@@
Finally, both $(\frac{\eta}{2}h-Ah)\overline{X}_{t}^{\prime 2}+\frac{\eta h}{2}\big(({\bar{x}_{0}^{N}})^{2}+2\overline{X}_{t}^{\prime}\bar{x}_{0}^{N}\big)-\eta\kappa\overline{X}_{t}^{\prime}$ and $\frac{\epsilon}{2}Q_{t}^{2}-\|C(t)\|_{\infty}|Q_{t}|$ have $O(1)$ lower bounds, so
$J_{soc}(\boldsymbol{v})\geq\frac{1}{N}\mathbb{E}\left[\sum_{i=1}^{N}\left(\int_{0}^{T}\left(\frac{\epsilon}{2}Q_{t}^{2}+\epsilon^{\prime}(v_{t}^{i})^{2}\right)dt\right)\right]+O(1),$
and the proof of Lemma 4.2 is complete.